OpenStudy (anonymous):

Use scientific notation to multiply the following two numbers. Express the answer using scientific notation; retain all nonzero decimal places. (1.6 * 10^4) * (9.9 * 10^6)

OpenStudy (anonymous):

1.6 x 9.9 = 15.84 10^4 x 10^6 = 10^10 So (1.6 x 10^4) x (9.9 x 10^6) = 15.84 x 10^10 = 1.584 = 10^11